Wilma's and Betty's pools both hold 10000 gallons of water. If Wilma's garden hose fills at a rate of 650 gallons per hour and Betty's at 600 gallons per hour, how much longer will it take Betty's pool to fill than Wilma's?
Answer by jorel1380(3719)

10000/650=15.384hrs.
10000/600=16.67hrs.
16.67-15.384=1.2827hrs difference